Six Data Tables (A, B, C, D, E, F) are shown below.

For each Table, decide which of two methods (table only  or   equation, graph, or table) can be used to represent the data.

For the purposes of this question, assume that we only know and can use two graph forms / equations:

  x= c t + d (straight line graph) and    x = b t2+ c t + d (quadratic curve graph).

A   

t x   
0 4
1 7
2 10
3 13
4 16

B

t x   
0 3
1 4
2 2
3 5
4 1

C.

t    x   
0 0
1 1
2 4
3 9
4 16

D.

t x   
0 0
1 10
2 4
3 19
4 6

E.

t x   
0 0
1 3
2 5
3 3
4 0

F.

t x   
0 10
1 6
2 2
3 -2
4 -6
Answer #1

A ut x=4,to Put の in O う×=Cヒ+ Put ヒ=,メ=7 ううこ C 4 うメニ 3tty Put (t,x), to), 3,/3, (1) - 3(2)+ 13=3(3) 16 3(4 AIl are trve hepreTany メ= (t+a Pot (t,x)0,3), (1,9 う3 う d= 3 C4-3= Pt Lt),2 2-2+3 s false Sowe メ= ctta Yejeut Taking x=bピナtta Rut (ちう- (9, 3),Foom t 7t+3 Put 3 false 5 rejet we メニbピ + Ctta (an use table only methoィ. 9 Takng メニbと+(も+d E)=Lo,0)),, Put Sロニd ニ btCta C0table only Similarly able only Euation, gah or table メニ 10-4ヒ